OpenStudy (anonymous):

MEDAL AND FOLLOW FOR CORRECT ANSWER. Q)How can 1/3 x − 2 = 1/4 x + 11 be set up as a system of equations? a)3y − x = −6 4y − x = 44 b)3y + x = −6 4y + x = 44 c)3y − 3x = −6 4y − 4x = 44 d)3y + 3x = −6 4y + 4x = 44

imqwerty (imqwerty):

1/3 x − 2 = 1/4 x + 11=y Equate the 1st nd 2nd eqs. separately with y nd simplify to get the 2 requirred equations.